3 Problem 17-4 Prior service cost; components of pension expense; present velue concepts [LO17-3,17-6] Sachs Brands defined benefit pension plan specifies annual retirement benefits equol lo 12% service years in al years sal y payable at the end of each year. Angela Davenport was hired by Sachs at the beginning of 2004 and is expected to retire at the end of 2038 after 35 years service. Her retirement is expected to span 18 years Davenports salary is $90,000 at the end of 2018 and the companys actuary projects her salary to be $280,000 at retirement. The actuarys discount rate is 8%。 Oft py St E o eBook PVA of S1, FVAD of $1 and PVAD of $) (Use appropriate factorís) from the tables provided.) Ask Prnt At the beginning of 2019, the pension formula was amended to: References 135% Service years Final years salary The amendment was made retroactive to apply the increased benefits to prior service years Required: 1. What is the companys prior service cost at the beginning of 2019 with respect to Davenport after the amendment described above? 2. Since the amendment occurred at the beginning of 2019, amortization of the prior service cost begins in 2019. What is the prior service cost amortization that would be included in pension expense? 3. What is the service cost for 2019 with respect to Davenport? 4. What is the interest cost for 2019 with respect to Davenport? 5. Calculate pension expense for 2019 with respect to Davenport, assuming plan assets attributable to her of $100,000 and a rate of return (actual and expected) of 10%. (For all requirements, do not round intermediate calculations. Round your final answers to nearest whole dollar) Next >
media%2Fc16%2Fc16c23cc-a39e-4ce2-aac5-f4
Answer #1
  1. what is the company prior service cost at the beginning of 2019 with respect to davenport after the amendment described above?

1.2%*14 years * $280,000 = 47070

47040* 9.3718 (n= 18, I =8%) = = 440850

440850*0.19865 = $87575

2.1.2%*1* $280000 = $3360

3. service cost for 2019 in respect of davenport

$3360*9.3718 (n= 18, I =8%) = $31490

$31,490* 0.21454 ( n= 20 i =8%) $6755

4. Interest cost for 2019 in respect of davenport

$87,575*8% = $7006